OpenStudy (anonymous):

When solving x2 – 20x = 5 by completing the square what value is added to both sides of the equation? –4 4 50 100

OpenStudy (anonymous):

x2 – 20x = 5 (x-10)^2 -100 = 5 -100 its -100

OpenStudy (anonymous):

because (x-10)^2 = x^2 -20 + 100